# False Sense of Security in Explainable Artificial Intelligence (XAI)

### v1

@@ Abstract
A cautious interpretation of AI regulations and policy in the EU and the USA place explainability as a central deliverable of compliant AI systems. However, from a technical perspective, explainable AI (XAI) remains an elusive and complex target where even state of the art methods often reach erroneous, misleading, and incomplete explanations. “Explainability” has multiple meanings which are often used interchangeably, and there are an even greater number of XAI methods – none of which presents a clear edge. Indeed, there are multiple failure modes for each XAI method, which require application-specific development and continuous evaluation. In this paper, we analyze legislative and policy developments in the United States and the European Union, such as the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence, the AI Act, the AI Liability Directive, and the General Data Protection Regulation (GDPR) from a right to explanation perspective. We argue that these AI regulations and current market conditions threaten effective AI governance and safety because the objective of trustworthy, accountable, and transparent AI is intrinsically linked to the questionable ability of AI operators to provide meaningful explanations. Unless governments explicitly tackle the issue of explainability through clear legislative and policy statements that take into account technical realities, AI governance risks becoming a vacuous “box-ticking” exercise where scientific standards are replaced with legalistic thresholds, providing only a false sense of security in XAI.
@@ 1 Introduction
Recent advancement in AI rely on end-to-end training of deep neural networks (DNNs), which act like black box algorithms. While superb performance of predictive and generative models led to their wide applications in all areas of society, from social media to healthcare, developers and users do not have a full grasp of their operating characteristics. The output (e.g., classification of an image) is obtained without understanding the AI’s decision making process. Yet, governance of AI demands revealing and reporting this mechanism as much as possible. On one hand, explanations are necessarily because fundamentally, humans need to understand automated decision making for meaningful governance. On the other hand, laws are being proposed and enacted to demand the AI companies to provide such explanations.
@@ 2 Legislative and Policy Developments
Generally, AI policy statements focus on making AI safe, secure, transparent, trustworthy, and explainable. Though these concepts receive less attention or detailed specification in the laws enacted to implement these policies. This led to ongoing debate about the extent to which a right to explanation exists. Based on relevant developments, we argue that some form of explainability will likely become a de facto requirement for critical AI systems deployed on the global market.
@@ 2.1 United States of America
In the United States, Joe Biden’s October 2023 Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence (14110) calls upon the executive agencies under his authority to study and to propose regulations that can effectively govern the safe and transparent use of AI in both the private and public sectors Biden (2023). National Institute of Standards and Technology (NIST), Department of Energy (DoE), National Science Foundation (NSF), and other agencies are crafting relevant resources, guidelines, and plans with industry and stakeholders.
@@
Enacted in 1974, the ECOA was not written with generative AI in mind, but rather the deterministic decision making systems in use at the time. Applied to contemporary technology, it nonetheless sets an actionable standard by ruling out the use of any black box AI systems that do not permit the isolation of the primary factors that contributed to their output. Such a clear line rule is effective and practicable, but only because the law applies to a narrow scope of activity. It remains to be seen whether the continuation of this piecemeal legislative approach to the use of algorithmic systems is desirable, as the multiplication of possible use cases of AI will almost certainly outpace the ability for the U.S. legislature to address them. At some point, broader regulations will be needed.
@@ 2.2.1 AI Act
The AI Act was endorsed by members of the European Parliament (MEP) in March 2024 European Parliament (2024), marking the first foray into the generalized regulation of AI. The Act and its relative merits have been discussed in detail elsewhere, and for the purposes of this paper it is sufficient to note that it does not promulgate any substantive and precise explainability thresholds for AI developers . In fact, the AI Act itself does explicitly not reference explainability, focusing instead on the need for transparency in high risk AI models.
@@
Transparency is addressed in Article 13(1) European Commission (2021), which states that “High-risk AI systems shall be designed and developed in such a way to ensure that their operation is sufficiently transparent to enable users to interpret the system’s output and use it appropriately. An appropriate type and degree of transparency shall be ensured.” The ethics guidelines of the European Commission (2019) clarify this requirement insofar as the decision making process must be recorded and documented, and an “understandable explanation of the algorithmic decision-making processes” must be published European Parliament (2019). According to Articles 6 and 7, high risk systems would include those deployed within the education, health, recruitment, critical infrastructure management, law enforcement, and justice domains because such use cases could result in threats to the health, safety, or fundamental rights of individuals European Commission (2021).
@@
As written, the AI Act does not provide any tangible thresholds for the transparency of high risk models. Instead, the specific technical requirements of the models affected by Article 13 will be defined through a collaborative approach by the governing institutions created by the Act: the AI Office, the European Artificial Intelligence Board, the Advisory Forum, and the Scientific Panel of Independent Experts. This decentralized approach is commendable because it represents a tacit admission from the European legislators that the issue of AI regulation is too complex for a single body, much less a political one, to take on alone. Only through a deliberative (and almost certainly iterative) process will specific thresholds be defined for high risk AI models.
@@
On the other hand, the silence on the issue of explainability is deafening. After several years of study and deliberation, the inability to seriously address the concept does not bode well for future efforts, possibly camouflaging its intractable issues. More precise yet possibly misaligned guidance from the AI Act may have been more useful than no guidance at all, as stakeholders across the globe are having to guess as to how their models may or may not meet future EU regulatory standards. One particular area of uncertainty is whether an expectation of explainability will apply to non high risk models in Europe. Such an assumption could be misguided, because a cautious reading of the GDPR and the AI Liability Directive shows that explainability will be – in some shape or form – a de facto compliance requirement for all AI models deployed in the EU.
@@ 2.2.3 AI Liability Directive
The final and potentially the most potent source of a right to explainability springs from the AI Liability Directive European Commission (2022). In this proposed directive, the European Parliament and Council have laid out a burden shifting scheme applying to fault-based liability claims for damages resulting from the output of AI systems. To avoid the harshest legal outcomes under this proposed burden shifting, deployers of AI models of all risk profiles should make their systems transparent and explainable.
Generally, establishing fault based liability requires the claimant to prove that the defendant’s fault has caused their damages. However, the AI Liability Directive would create under some conditions a presumption that a claimant’s damages have been caused by the defect or fault in the AI system deployed by the defendant. This effectively shifts the burden to the defendant to prove that their system did not cause the claimed damages. This policy is an acknowledgment that placing the burden of proving causation on the claimant would drastically reduce their chances of recovery in almost all cases because of the inherent difficulty in “connecting the dots” between an AI system’s inner workings and its outputs.
The presumption of causation always applies to claims against high risk AI systems, except “where the defendant demonstrates that sufficient evidence and expertise is reasonably accessible for the claimant to prove the causal link” European Commission (2022)(art 4.4). To limit the risk of liability, the deployers of high risk models should ensure that they are not using black box models, and publish documentation and other assets that adequately explain the high risk model’s behavior. Doing so would presumably shift the burden of causation back to potential claimants.
For non-high risk models, the presumption of causality would only apply where, according to Article 4.5, “the national court considers it excessively difficult for the claimant to prove the causal link” European Commission (2022). Since there is no consensus as to what would be a scientifically or legally sufficient threshold of explaining the outputs of a generative AI system, it’s unclear if or how anyone could establish with clear and convincing evidence that a fault in an AI system has proximately caused its harmful outputs. As a result, national courts in practice might interpret this section (as enacted locally) in a manner more sympathetic to claimants than intended by the EU legislature. This is a key area where again, deployers of non-high risk models can avoid potential liability by pro-actively taking steps to make their systems explainable.
@@ 3 Polysemy of “Explainability”
The term “explainability” is used in a wide array of scenarios. Similarly, trustworthiness, safety, and related terms are used, often without clear guidelines. While there have been efforts to taxonomize qualities, formats, data types, and purposes of XAI, those conceptual definitions are difficult to apply and may not clearly delineate technical development in practice. Especially seeing from how legistrative texts are written, “explainability” has become a catch-all phrase, which are up for broad interpretation. For comprehensive reviews and taxonomical organizations, see Stepin et al. (2021); Barredo Arrieta et al. (2020); Nauta et al. (2023); Saeed and Omlin (2023). Here we highlight five areas that any explainability regulation should address specifically:
First, there are two approaches to explaining AI; global and local. Global explanations attempt to answer how the AI model works, as an overall characterization obtained before a specific sample has been processed,ex ante. Furthermore, global explanations look at overall relationships between variables, which are especially critical when dealing with non-linear relationships. Local explanations, on the other hand, refer to how the AI model worked on a particular sample post hoc. Often, a non-linear relationship can be characterized linearly in a small window (local) of some features. Or, in classifying and predicting based on a specific sample image, the model may have used a small subset of features.
@@
Third, generally speaking, the current XAI methodology can’t give comprehensive yet human-understanable explanations for blackbox AI. Except in exceptional scenarios, only the full model and weights can fully describe the “complete mechanisms” of deep neural networks. In general, when explanations are complete, they could simply be used as an efficient and interpretable model with equivalent performance. In practice, post-hoc explainability methods are substantially reducing the model mechanisms and characterizing different aspects of the model, e.g., by a number of competing importance estimators (also known as saliency maps).
@@
Practically, the explainability requirements of consumer protection regulations such as the U.S. Equal Credit Opportunity Act U.S. Department of Justice (1974) and the GDPR European Parliament and Council of the European Union (2016b) imply that the provided explanation should be understandable by the average person. On the other hand, explanations published by corporations hoping to avoid the presumption of causality promulgated under the EU AI Liability Directive European Commission (2022) would need to be much more detailed. Overcoming a legally-mandated presumption relating to causality cannot be achieved through pedestrian means, likely requiring the publication of data sets and methodologies that illuminate exactly how the claimant’s damages were (not) caused by the AI system’s alleged fault.
Fifth, being able to explain the model’s decision making process is related to the ability to control and interact with the AI model in a meaningful way Teso and Kersting (2019)Schmid and Finzel (2020). While explanations may (truthfully) characterize the model, how to change its behavior according to our desire is not always feasible. One may correct a misbehavior (e.g., mis-classification) by re-labeling data, re-training the model, and even including certain pre- and post-processing steps, but they remain an unsolved challenge. We may encourage the model to work closer to how we want it to behave, but short of compiling hard-coded rules [^1] Rule-based algorithms exist and are useful in a wide range of areas. Those algorithms are not considered AI in this position paper., there’s no guarantee these guardrails will be effective.
The aforementioned five aspects of explainability are closely related. For example, local explanations may be reasonable and faithful for a given sample; however they may not provide sufficient information to control the model. Furthermore, the simplification of AI’s decision making processes, which is necessary for most complex non-linear models (e.g., DNNs), must be informed by the target audience. Beyond these five categories, there exist other definitions of explainability. Mechanistic explanations focuses on parts of the model (e.g., layers, sub-networks), whereas functional explanations are about the model’s functions Páez (2019). Ginsberg (1986); Grahne (1998) examine counterfactual explanations which can be used to explain black box models. Overall, these different interpretations of explainability would significantly impact what must be explained, assuming the technical challenges are solved. Even with the best intentions and improved technical capabilities, misalignment between the intention of laws and the implementation of AI companies will create confusing and uninformative processes.
@@ 4.0.1 Robustness
First, XAI methods may not be robust to small changes in inputs Alvarez-Melis and Jaakkola (2018). For example, in Figure 2, three samples are based on the same image of ‘3’. For this experiment, small and independent noise is added. A simple neural network with one convolutional layer is trained on the MNIST database Deng (2012). The trained CNN model correctly classifies all three images with high confidence. Nonetheless, explanations by Local Interpretable Model-agnostic Explanations (LIME; on the second row) Ribeiro et al. (2016) are substantially varied. The positive and negative influence of pixels, as estimated Figure 2, would present drastically different interpretation on how the classification has been made.
Relatedly, Adebayo et al. (2018) investigated robustness of saliency maps, which are designed to quantify and highlight important pixels with respect to the class label. They demonstrated that many advanced saliency maps show similar explanations even with data randomization. In the worst cases, some saliency maps, which are supposed to provide local explanations based on the model, resemble edge detection, which are by design independent of the model or the classification Adebayo et al. (2018). Generally, these types of problems may be mitigated by optimizing for robustness Alvarez-Melis and Jaakkola (2018) or selecting appropriate methods Adebayo et al. (2018).
@@ 4.0.2 Adversarial Attacks
Second, adversarial attacks on XAI have been investigated, such that even well-performing methods give nonsensical explanations Dombrowski et al. (2019). Instead of independent noise added in the first failure mode, the attackers can specifically design imperceptible changes to the input (e.g., the image looks identical to human before and after modification), which completely change the explanations. This type of vulnerability mirrors a well-known adversarial attacks on DNNs where carefully designed imperceptible changes to the input can cause the model to misbehave and misclassify. Mitigation strategies for adversarial attacks include explanation aggregation and model regularization Baniecki and Biecek (2023). Nonetheless, attacks and defenses surrounding explanations will continue to evolve as XAI become more mainstream and public-facing.

### v2

@@ Abstract
A cautious interpretation of AI regulations and policy in the EU and the USA places explainability as a central deliverable of compliant AI systems. However, from a technical perspective, explainable AI (XAI) remains an elusive and complex target where even state-of-the-art methods often reach erroneous, misleading, and incomplete explanations. “Explainability” has multiple meanings which are often used interchangeably, and there are an even greater number of XAI methods – none of which presents a clear edge. Indeed, there are multiple failure modes for each XAI method, which require application-specific development and continuous evaluation. In this paper, we analyze legislative and policy developments in the United States and the European Union, such as the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence, the AI Act, the AI Liability Directive, and the General Data Protection Regulation (GDPR) from a right to explanation perspective. We argue that these AI regulations and current market conditions threaten effective AI governance and safety because the objective of trustworthy, accountable, and transparent AI is intrinsically linked to the questionable ability of AI operators to provide meaningful explanations. Unless governments explicitly tackle the issue of explainability through clear legislative and policy statements that take into account technical realities, AI governance risks becoming a vacuous “box-ticking” exercise where scientific standards are replaced with legalistic thresholds, providing only a false sense of security in XAI.
@@ 1 Introduction
Recent advancements in AI rely on end-to-end training of deep neural networks (DNNs), which act like black box algorithms. While the superb performance of predictive and generative models led to their wide applications in all areas of society, from social media to healthcare, developers, and users do not have a full grasp of their operating characteristics. The output (e.g., classification of an image) is obtained without understanding the AI’s decision-making process. Yet, governance of AI demands revealing and reporting this mechanism as much as possible. Explanations are necessary because fundamentally, humans need to understand automated decision-making for meaningful governance. Furthermore, laws are being proposed and enacted to demand that AI companies provide such explanations.
@@ 2 Legislative and Policy Developments
Generally, AI policy statements focus on making AI safe, secure, transparent, trustworthy, and explainable. These concepts, however, receive less attention or detailed specification in the laws enacted to implement these policies. This led to ongoing debate about the extent to which a right to explanation exists. Based on relevant developments, we argue that some form of explainability will likely become a de facto requirement for critical AI systems deployed on the global market.
@@ 2.1 United States of America
In the United States, Joe Biden’s October 2023 Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence (14110) calls upon the executive agencies under his authority to study and propose regulations that can effectively govern the safe and transparent use of AI in both the private and public sectors Biden (2023). The National Institute of Standards and Technology (NIST), Department of Energy (DoE), National Science Foundation (NSF), and other agencies are crafting relevant resources, guidelines, and plans with industry and stakeholders.
@@
Enacted in 1974, the ECOA was not written with generative AI in mind, but rather the deterministic decision-making systems in use at the time. Applied to contemporary technology, it nonetheless sets an actionable standard by ruling out the use of any black box AI systems that do not permit the isolation of the primary factors that contributed to their output. Such a clear-line rule is effective and practicable, but only because the law applies to a narrow scope of activity. It remains to be seen whether the continuation of this piecemeal legislative approach to the use of algorithmic systems is desirable, as the multiplication of possible use cases of AI will almost certainly outpace the ability of the U.S. legislature to address them. At some point, broader regulations will be needed.
@@ 2.2.1 AI Act
The AI Act was endorsed by members of the European Parliament (MEP) in March 2024 European Parliament (2024), marking the first foray into the generalized regulation of AI. The Act and its relative merits have been discussed in detail elsewhere, and for the purposes of this paper, it is sufficient to note that it does not promulgate any substantive and precise explainability thresholds for AI developers. In fact, the AI Act itself does explicitly not reference explainability, focusing instead on the need for transparency in high-risk AI models.
@@
Transparency is addressed in Article 13(1) European Commission (2021), which states that “High-risk AI systems shall be designed and developed in such a way to ensure that their operation is sufficiently transparent to enable users to interpret the system’s output and use it appropriately. An appropriate type and degree of transparency shall be ensured.” The ethics guidelines of the European Commission (2019) clarify this requirement insofar as the decision-making process must be recorded and documented, and an “understandable explanation of the algorithmic decision-making processes” must be published European Parliament (2019). According to Articles 6 and 7, high-risk systems would include those deployed within the education, health, recruitment, critical infrastructure management, law enforcement, and justice domains because such use cases could result in threats to the health, safety, or fundamental rights of individuals European Commission (2021).
@@
As written, the AI Act does not provide any tangible thresholds for the transparency of high-risk models. Instead, the specific technical requirements of the models affected by Article 13 will be defined through a collaborative approach by the governing institutions created by the Act: the AI Office, the European Artificial Intelligence Board, the Advisory Forum, and the Scientific Panel of Independent Experts. This decentralized approach is commendable because it represents a tacit admission from the European legislators that the issue of AI regulation is too complex for a single body, much less a political one, to take on alone. Only through a deliberative (and almost certainly iterative) process will specific thresholds be defined for high-risk AI models.
@@
On the other hand, the silence on the issue of explainability is deafening. After several years of study and deliberation, the inability to seriously address the concept does not bode well for future efforts, possibly camouflaging its intractable issues. More precise yet possibly misaligned guidance from the AI Act may have been more useful than no guidance at all, as stakeholders across the globe are having to guess as to how their models may or may not meet future EU regulatory standards. One particular area of uncertainty is whether an expectation of explainability will not apply to non-high-risk models in Europe. Such an assumption could be misguided, because a cautious reading of the GDPR and the AI Liability Directive shows that explainability will be – in some shape or form – a de facto compliance requirement for all AI models deployed in the EU.
@@ 2.2.3 AI Liability Directive
The final and potentially the most potent source of a right to explainability springs from the AI Liability Directive European Commission (2022). In this proposed directive, the European Parliament and Council have laid out a burden-shifting scheme applying to fault-based liability claims for damages resulting from the output of AI systems. To avoid the harshest legal outcomes under this proposed burden-shifting, deployers of AI models of all risk profiles should make their systems transparent and explainable.
Generally, establishing fault-based liability requires the claimant to prove that the defendant’s fault has caused their damages. However, the AI Liability Directive would create under some conditions a presumption that a claimant’s damages have been caused by the defect or fault in the AI system deployed by the defendant. This effectively shifts the burden to the defendant to prove that their system did not cause the claimed damages. This policy is an acknowledgment that placing the burden of proving causation on the claimant would drastically reduce their chances of recovery in almost all cases because of the inherent difficulty in “connecting the dots” between an AI system’s inner workings and its outputs.
The presumption of causation always applies to claims against high-risk AI systems, except “where the defendant demonstrates that sufficient evidence and expertise is reasonably accessible for the claimant to prove the causal link” European Commission (2022)(art 4.4). To limit the risk of liability, the deployers of high-risk models should ensure that they are not using black box models, and publish documentation and other assets that adequately explain the high-risk model’s behavior. Doing so would presumably shift the burden of causation back to potential claimants.
For non-high-risk models, the presumption of causality would only apply where, according to Article 4.5, “the national court considers it excessively difficult for the claimant to prove the causal link” European Commission (2022). Since there is no consensus as to what would be a scientifically or legally sufficient threshold of explaining the outputs of a generative AI system, it’s unclear if or how anyone could establish with clear and convincing evidence that a fault in an AI system has proximately caused its harmful outputs. As a result, national courts in practice might interpret this section (as enacted locally) in a manner more sympathetic to claimants than intended by the EU legislature. This is a key area where again, deployers of non-high risk models can avoid potential liability by pro-actively taking steps to make their systems explainable.
@@ 3 Polysemy of “Explainability”
The term “explainability” is used in a wide array of scenarios. Similarly, trustworthiness, safety, and related terms are used, often without clear guidelines. While there have been efforts to taxonomize qualities, formats, data types, and purposes of XAI, those conceptual definitions are difficult to apply and may not clearly delineate technical development in practice. Especially seeing from how legislative texts are written, “explainability” has become a catch-all phrase, which is up for broad interpretation. For comprehensive reviews and taxonomical organizations, see Stepin et al. (2021); Barredo Arrieta et al. (2020); Nauta et al. (2023); Saeed and Omlin (2023). Here we highlight five areas that any explainability regulation should address specifically:
First, there are two approaches to explaining AI; global and local. Global explanations attempt to answer how the AI model works, as an overall characterization obtained before a specific sample has been processed, ex ante. Furthermore, global explanations look at overall relationships between variables, which are especially critical when dealing with non-linear relationships. Local explanations, on the other hand, refer to how the AI model worked on a particular sample post hoc. Often, a non-linear relationship can be characterized linearly in a small window (local) of some feature values. Or, in classifying and predicting based on a specific sample image, the model may have used a small subset of features.
@@
Third, generally speaking, the current XAI methodology can’t give comprehensive yet human-understandable explanations for black box AI. Only the full model and weights can truly describe the “complete mechanisms” of deep neural networks. In general, when explanations are complete, they could simply be used as an efficient and interpretable model with equivalent performance. In practice, post-hoc explainability methods substantially reduce the model mechanisms and characterize different aspects of the model, e.g., by estimating and visualizing the relative importance of input features (also known as saliency maps).
@@
Practically, the explainability requirements of consumer protection regulations such as the U.S. Equal Credit Opportunity Act U.S. Department of Justice (1974) and the GDPR European Parliament and Council of the European Union (2016b) imply that the provided explanation should be understandable by the average person. On the other hand, explanations published by corporations hoping to avoid the presumption of causality promulgated under the EU AI Liability Directive European Commission (2022) would need to be much more detailed. Overcoming a legally mandated presumption relating to causality cannot be achieved through pedestrian means, likely requiring the publication of data sets and methodologies that illuminate exactly how the claimant’s damages were (not) caused by the AI system’s alleged fault.
Fifth, being able to explain the model’s decision-making process is related to the ability to control and interact with the AI model in a meaningful way Teso and Kersting (2019)Schmid and Finzel (2020). While explanations may (truthfully) characterize the model, how to change its behavior according to our desire is not always feasible. One may correct a misbehavior (e.g., misclassification) by re-labeling data, re-training the model, and even including certain pre- and post-processing steps, but this remains an unsolved challenge. We may encourage the model to work closer to how we want it to behave, but short of compiling hard-coded rules [^1] Rule-based algorithms exist and are useful in a wide range of areas. Those algorithms are not considered AI in this position paper., there’s no guarantee these guardrails will be effective.
The aforementioned five aspects of explainability are closely related. For example, local explanations may be reasonable and faithful for a given sample; however, they may not provide sufficient information to control the model. Furthermore, the simplification of AI’s decision-making processes, which is necessary for most complex non-linear models (e.g., DNNs), must be informed by the target audience. Beyond these five categories, there exist other definitions of explainability. Mechanistic explanations focus on parts of the model (e.g., layers, sub-networks), whereas functional explanations are about the model’s functions Páez (2019). Ginsberg (1986); Grahne (1998) examine counterfactual explanations which can be used to explain black box models. Overall, these different interpretations of explainability would significantly impact what must be explained, assuming the technical challenges are solved. Even with the best intentions and improved technical capabilities, misalignment between the intention of laws and the implementation of AI companies will create confusing and uninformative processes.
@@ 4.0.1 Robustness
First, XAI methods may not be robust to small changes in inputs Alvarez-Melis and Jaakkola (2018). For example, in Figure 2, three samples are based on the same image of ‘3’. For this experiment, small and independent noise is added. A simple neural network with one convolutional layer is trained on the MNIST database Deng (2012). The trained CNN model correctly classifies all three images with high confidence. Nonetheless, explanations by Local Interpretable Model-agnostic Explanations (LIME; on the second row) Ribeiro et al. (2016) are substantially varied. The positive and negative influence of pixels, as estimated Figure 2, would present drastically different interpretations of how the classification has been made.
Relatedly, Adebayo et al. (2018) investigated the robustness of saliency maps, which are designed to quantify and highlight important pixels with respect to the class label. They demonstrated that many advanced saliency maps show similar explanations even with data randomization. In the worst cases, some saliency maps, which are supposed to provide local explanations based on the model, resemble edge detection, which is by design independent of the model or the classification Adebayo et al. (2018). Generally, these types of problems may be mitigated by optimizing for robustness Alvarez-Melis and Jaakkola (2018) or selecting appropriate methods Adebayo et al. (2018).
@@ 4.0.2 Adversarial Attacks
Second, adversarial attacks on XAI have been investigated, such that even well-performing methods give nonsensical explanations Dombrowski et al. (2019). Instead of independent noise added in the first failure mode, the attackers can specifically design imperceptible changes to the input (e.g., the image looks identical to a human before and after modification), which completely change the explanations. This type of vulnerability mirrors well-known adversarial attacks on DNNs where carefully designed imperceptible changes to the input can cause the model to misbehave and misclassify. Mitigation strategies for adversarial attacks include explanation aggregation and model regularization Baniecki and Biecek (2023). Nonetheless, attacks and defenses surrounding explanations will continue to evolve as XAI becomes more mainstream and public-facing.